Summary: Sky was clear initially, but clouds started to move in around 5:30 UT. Seeing was sub-arcsecond at the beginning, but quickly deteriorated to 1.5"-2". The dome was closed around 8:00 UT due to thickening clouds; It was reopened at 11:00 UT. We executed two OBs for Olszewski's program and 1.5 OBs for Jiang's program. We also executed two backup OBs for Milne's program, but these are probably affected by thin patchy clouds and a bad seeing. One OB was also executed to observe a standard-star field. A frustrating night.
  • Problem - The slow connection problem to this wiki page persisted. Connection was often (but not always) very slow (especially late at night), sometimes taking a few to several minutes to save/open the log. In the worst case, attempts to save/open the log would fail with the "Internal Server Error" after a long wait. This problem seems specific to this wiki page since there was no problem with accessing other sites. A temporary solution may be to keep a log in a separate file and upload it at the end of the night.


1:50
  • Opening the dome. Clear.
2:00
  • Start taking twilight flats: blue-g, r; red-r, F970N20
2:35
  • Pointing check near zenith.
2:50
  • Focusing/collimating on a star close to the first target.
3:15
  • UMaIgr_pos0 (E. Olszewski) started.
    Estimated seeing with dofpia: ~1" in red, ~1.8" in blue.
    Measured seeing in the 1st images: ~0.9" in red, ~1.5" in blue.
    Images look much rounder in the Blue channel today.
    ... well, image quality deteriorated quickly after the 1st frames.
    OB paused after third images, focused/collimated, and OB restarted at 3:35.
    Measured seeing: ~0.7" in red, ~0.8" in blue (but elongated).
    Focussing/collimating.
4:00
  • UMaIgr_pos1 (E. Olszewski) started.
4:35
  • Moving to SA104 standard star field.
    Focusing/collimating.
4:45
  • SA104_BgVrrIN (SA104 multi-filter observations) started.
    B, g, V, r, r, I, F970N20.
4:55
  • Going to a pointing star close to the next target. Focusing/collimating. Adjusting the pointing model.
    Going to the next target (IRAC dark field). Focusing/collimating.
5:10
  • IRACdark_grN8 (L. Jiang) started.
    Clouds started to come in 5:25...
    OB aborted at 5:30.
5:40
  • Backup OB Milne_SN06E (P. Milne) started after focus/collimation.
    Looking through clouds; Seeing getting worse (~1.5" in blue).
6:10
  • Backup OB Milne_SN09dc_rI (P. Milne) started after focus/collimation.
    Sky gradually clearing up. Often encountered "Internal Server Error" when trying to save the log.
6:40
  • IRACdark_grN8 (L. Jiang) started.
    OB aborted because of a bad seeing (~2").
    focusing/collimating.
    Estimated seeing with dofpia: 1.6" in blue, 1.5" in red.
6:55
  • Going to a nearby star for a pointing.
    Moving back to the science target. Focusing/collimating.
7:10
  • IRACdark_grN8 (L. Jiang) started.
    Clouds coming in again... Focusing/Collimating, but difficult due to decreasing signal.
8:00
  • Completely covered up by clouds...
    Dome closed.
11:00
  • Sky clearing up. Opening the dome again.
11:15
  • IRACdark_grN8 (L. Jiang) started after focus/collimation.
    Good seeing ~0.7"; sky background ~ 9000 DN.
11:50
  • IRACdark_grN9 (L. Jiang) started after focus/collimation.
    OB length reduced due to the morning twilight.
12:15
  • Taking twilight flats for Milne's program: B, V, r, I

-- EiichiEgami - 16 Apr 2010
Topic revision: r3 - 16 Apr 2010, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback